# Fully heavy tetraquark states with diquark-antidiquark configuration

### v1

@@ Abstract
The study of the fully heavy tetraquark state helps to better understand the characteristics of compact exotic hadrons formed by quarks and gluons. In this work, we systematically investigate the mass spectra and decays of fully heavy tetraquark states using a diquark-antidiquark model based on the one gluon exchange process. The tetraquark states are treated as compact diquark–antidiquark configurations, denoted as $QQ][\bar{Q}\bar{Q}]$, where $Q=c$ or $b$. By properly accounting for the Pauli principle, basis wave functions incorporating flavor, color, and spin are constructed. The required model parameters are derived from fitting conventional hadron experimental values and lattice QCD calculations. We calculate the mass spectra and decays of the fully heavy tetraquarks in the lower $S$- and $P$-waves. Our results do not support interpreting $X(6600)$, $X(6900)$, and $X(7200)$ as $S$-wave fully charmed tetraquark states. Nevertheless, our analysis suggests that $X(6600)$ can be a good candidate for the $P$-wave fully charmed tetraquark state. Furthermore, we also identifies several possible fully heavy tetraquark states that may be observable in future experiments.
@@ I Introduction
Quantum Chromodynamics (QCD) allows for a variety of exotic hadronic structures beyond conventional mesons and baryons, such as multiquark states, glueballs, and hybrid states [1, 2, 3]. Among these, tetraquark states have become a major focus in hadron physics. Since the Belle Collaboration’s discovery of the $X(3872)$ resonance in 2003 [4], a growing number of exotic hadrons have been observed experimentally [5]. These discoveries not only confirm the existence of exotic hadronic states but also inspire further theoretical and experimental investigations into their properties. Among the tetraquark states, fully heavy tetraquarks represent a particularly intriguing class. In such systems, light-meson exchange is forbidden, and the interaction is dominated by short-range color forces mediated by gluon exchange, suggesting a compact multiquark configuration.
@@
In recent years, several candidates for fully charmed tetraquark states have been observed experimentally. In 2020, the LHCb Collaboration observed a broad structure above twice the $J/\psi$ mass threshold ranging from 6.2 to 6.8 GeV and a narrower structure at about 6.9 GeV, called $X(6900)$ [6]. Very recently, the CMS and ATLAS Collaborations also researched the di-$J/\psi$ channel and both of them confirmed the existence of the $X(6900)$ [7, 8]. Besides, the CMS Collaboration reported two new structures, the $X(6600)$ and $X(7200)$, in the di-$J/\psi$ invariant mass spectrum [7]. The ATLAS Collaboration also observed the new structures, $X(6400)$ and $X(6600)$, in the di-$J/\psi$ channel and $X(7200)$ in the $J/\psi$-$\psi(2S)$ invariant mass spectrum [8]. These experiments have motivated a large number of theoretical studies on fully charmed tetraquark states [9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Many theoretical studies suggest that $X(6900)$ is a $2S$-wave tetraquark state [23, 24, 25, 26, 27, 28, 16], while others propose it as an $S$-wave state with angular quantum number $L=0$ and total quark spin $S=2$ [29]. Some works interpret $X(6600)$ as a $2S$-wave state [28, 30], whereas others regard it as a $1S$-wave [23]. There are also studies that suggest $X(7200)$ could be a $3S$-wave [27, 28] or a $2P$-wave state [16].
@@
Furthermore, some new experimental evidence for fully bottom tetraquark states has been reported. In 2016, the CMS Collaboration announced the first observation of $\Upsilon(1S)$ pair production in $pp$ collisions at $\sqrt{s}=8$ TeV [31]. They found an excess with a global significance of 3.6 $\sigma$ in the $4\mu$ channel. The invariant mass of this excess is $18.4\pm 0.1\pm 0.2$ GeV [32]. Two years later, the LHCb Collaboration performed a research for the decay of the beautiful tetraquark, $X$, to the $\Upsilon(1S)\mu^{+}\mu^{-}$ final state. However, no significant excess is seen for any mass hypothesis in the range [17.5, 20.0][17.5,\ 20.0] GeV [33]. In 2019, other production of double $\Upsilon(1S)$ is observed in Cu + Au collisions at the Relativistic Heavy Ion Collider by ANDY Collaboration. They shows evidence of a significant peak at $M=18.12\pm 0.15\pm 0.6$ GeV [34], which is compatible with the first observation of an all-$b$ tetraquark. The fully bottom tetraquark states have also been the subject of extensive theoretical investigations [9, 35, 36, 37, 20]. Given that only a few experiments have observed these possible fully bottom tetraquark states, further experimental investigations are required to confirm their existence.
@@
Fully heavy tetraquark states, especially the fully charmed and fully bottom tetraquark states, have attracted extensive attention from theorists. A wide range of theoretical approaches have been employed to study them, including the QCD sum rules [38, 10, 39, 11], a variety of quark models [24, 13, 16, 40, 18, 20, 41, 42, 9, 43], the diquark models [12, 35, 19, 25], the chromomagnetic model [44, 45, 46], the diffusion Monte Carlo approach [36, 47, 48], and the other phenomenological models [26, 27, 28, 29, 30]. A few studies suggest that the fully heavy tetraquark states $cc\bar{c}\bar{c}$ and $bb\bar{b}\bar{b}$ could be bound states. In Ref. [46], the lowest mass of the $cc\bar{c}\bar{c}$ system lies below the di-$\eta_{c}$ threshold, while the lightest S-wave $bb\bar{b}\bar{b}$ state appears below the $\eta_{b}\eta_{b}$ threshold in Ref. [35, 49, 9]. However, the majority of investigations conclude that the $S$-wave fully heavy tetraquarks are located above the corresponding double-meson thresholds. For example, in Ref. [50], the authors used the first-principles lattice nonrelativistic QCD methodology to study the low-lying spectrum of the $bb\bar{b}\bar{b}$ tetraquark states with quantum numbers $0^{++}$, $1^{+-}$ and $2^{++}$. However, they found no evidence of a QCD bound tetraquark below the lowest non-interacting thresholds in the channels studied. It is evident that theorists hold diverse views regarding the new experimental observations, offering a variety of interpretations and even arriving at some contradictory conclusions. At present, there is no theoretical explanation that has gained universal acceptance. The search for new theoretical approaches remains an active and ongoing endeavor.
@@
Nowadays, a variety of new exotic hadrons are continuously being observed in experiments, while theoretical progress remains relatively limited. Further theoretical studies are still required to deepen our understanding of the mechanisms underlying strong interactions. In this work, we investigate the mass spectra of fully heavy tetraquark states within the diquark–antidiquark configuration and analyze their strong decay behaviors.
@@
This paper is organized as follows. In Sec. II, we present the effective Hamiltonian employed in this work and provide the basis wave functions corresponding to different quantum numbers. Subsequently, we introduce the formulae used for calculating the strong decay widths. In Sec. III, we obtain the model parameters required for the calculations and describe the fitting methods used to obtain these parameters. We then show our calculated results and provide an analysis of the decay processes. We summarize in Sec. IV. The nature units $c=\hbar=1$ are used throughout.
@@ II.1 Effective Hamiltonian
The bound state system formed by four fully heavy quarks can be seen as a unique example of direct interaction between quarks and gluons. Consequently, the interaction in this system is mainly manifested as the one gluon exchange process [51, 52]. For the fully heavy tetraquark systems, the improved nonrelativistic effective Hamiltonian reads [53, 54]
@@
$H_{SS}={}$ $\sum_{i<j}2\kappa_{ij}(\bm{S}_{i}\cdot\bm{S}_{j}),$ (2)
@@
$H_{L}={}$ $B\frac{L(L+1)}{2}.$
Here, $m_{i}$ denotes the mass of the $i$-th constituent quark mass. $H_{SS}$ represents the interaction of each quark pair in the tetraquark, $H_{SL}$ and $H_{L}$ describe spin-orbit coupling and pure orbital contribution, respectively. $\bm{S}_{\mathcal{Q}}$ in $H_{SL}$ represents the spin operator of the diquark configuration $\mathcal{Q}=[Q_{1}Q_{2}]$. The orbital angular momentum $\bm{L}$ describes the relative motion between the diquark $\mathcal{Q}$ and the anti-diquark $\bar{\mathcal{Q}}^{\prime}$. The coefficients $A$ and $B$ are phenomenological parameters to be determined from experimental data.
@@ II.2 Basis wave functions
To obtain the mass spectra of the fully heavy tetraquark states, we need to determine their wave functions. Here we only show the wave functions of ground states and we represent them as the form $|{(Q_{1}Q_{2})_{s}^{c}(\bar{Q}_{3}\bar{Q}_{4})_{s}^{c}}\rangle$, where the superscript denotes the color and the subscript denotes the spin.
@@
The wave functions of the ground states are as follows, (1) $J^{P}=0^{+}$
@@
(2) $J^{P}=1^{+}$
@@
(3) $J^{P}=2^{+}$
@@
The $\bar{\delta}_{12}$ ($\bar{\delta}_{34}$) means that if $Q_{1}Q_{2}$ ($\bar{Q}_{3}\bar{Q}_{4}$) have the same flavor then this state can not exist. We also consider the charge conjugation and excited states with $L=1$. The complete wave functions are listed in Appendix A.
@@ II.3 Decay width
By inserting the basis vectors of the wave function into the Hamiltonian, one can obtain the mass spectra of the tetraquark states. We can then calculate their decays into different meson final states.
For a particle of mass $M$ and four-momentum $P$ decaying into two particles with momenta $p_{1}$ and $p_{2}$, the partial decay rate is written as
$d\Gamma=\frac{(2\pi)^{4}}{2M}|\mathcal{A}|^{2}d\Phi(P;p_{1},p_{2}),$ (7)
where $\mathcal{A}$ is the amplitude connecting the initial state to the final state and $d\Phi$ is an element of 2-body phase space given by
$d\Phi(P;p_{1},p_{2})$ $=\delta^{4}(P-p_{1}-p_{2})$ (8)
$\times\frac{d^{3}p_{1}}{(2\pi)^{3}2E_{1}}\frac{d^{3}p_{2}}{(2\pi)^{3}2E_{2}}.$
@@
$|\bm{p}_{1}|=|\bm{p}_{2}|=\frac{1}{2M}\sqrt{\lambda(M^{2},m_{1}^{2},m_{2}^{2})}$ (9)
where $\lambda(\alpha,\beta,\gamma)=\alpha^{2}+\beta^{2}+\gamma^{2}-2\alpha\beta-2\alpha\gamma-2\beta\gamma$ is the Källén function. Thus, the partial decay rate can be written as
@@ III.1 Model parameters
To compute the ground-state masses of fully heavy tetraquark states, we need to determine the constituent masses of the charm and bottom quarks, as well as the parameters $\kappa_{ij}$ associated with quark pairs. We obtain the constituent masses and the parameters $\kappa_{ij}$ from known conventional hadrons. For conventional hadrons that have not yet been observed experimentally, we use data from Lattice QCD [55, 56].
The color states associated with the interaction parameters $\kappa_{ij}$ obtained through conventional hadron spectrum must exclusively correspond to either 1 or $\bar{3}$. Thus we first determine $(\kappa_{Q\bar{Q}})_{1}$ with mesons and $(\kappa_{QQ})_{\bar{3}}$ with baryons.
@@
$M=m_{Q_{1}}+m_{Q_{2}}+(\kappa_{Q_{1}\bar{Q}_{2}})_{1}[J(J+1)-\frac{3}{2}].$ (12)
With these equation we can calculate the value of $(\kappa_{Q_{1}\bar{Q}_{2}})_{1}$. Their numerical results are tabulated in Table 1.
Table 1: Parameter $\kappa$ (in MeV) for $q\bar{q}$ with color configuration of singlet and for qqqq with color anti-triplet. And these parameters are derived from fitting the $S$-wave mesons and baryons.
$(\kappa_{c\bar{c}})_{1}$ $(\kappa_{b\bar{b}})_{1}$ $(\kappa_{b\bar{c}})_{1}$ $(\kappa_{cc})_{\bar{3}}$ $(\kappa_{bb})_{\bar{3}}$ $(\kappa_{bc})_{\bar{3}}$
@@
To determine $(\kappa_{QQ})_{\bar{3}}$ with baryons, Eq. (1) can be reformulated as
$M=$ $\sum_{i=1}^{3}m_{i}+2\kappa_{q_{1}q_{2}}(\bm{s}_{q_{1}}\cdot\bm{s}_{q_{2}})+\kappa_{q_{1}q_{3}}(\bm{s}_{q_{1}}\cdot\bm{s}_{q_{3}})$
$+$ $\kappa_{q_{2}q_{3}}(\bm{s}_{q_{2}}\cdot\bm{s}_{q_{3}}).$ (13)

### v2

@@ Abstract
The study of the fully heavy tetraquark state helps to better understand the characteristics of compact exotic hadrons formed by quarks and gluons. In this work, we systematically investigate the mass spectra and decay channels of fully heavy tetraquark states using a diquark-antidiquark model based on the one gluon exchange process. The tetraquark states are treated as compact diquark–antidiquark configurations, denoted as $QQ][\bar{Q}\bar{Q}]$, where $Q=c$ or $b$. By properly accounting for the Pauli principle, basis wave functions incorporating flavor, color, and spin are constructed. The required model parameters are derived from fitting conventional hadron experimental values and lattice QCD calculations. We calculate the mass spectra and decays of the fully heavy tetraquarks in the lower $S$- and $P$-waves. Our results do not support interpreting $X(6600)$, $X(6900)$, and $X(7200)$ as the $1S$-wave fully charmed tetraquark states. Nevertheless, our analysis suggests that $X(6200)$ can be a good candidate for the $1S$-wave fully charmed tetraquark state with quantum numbers $2^{++}$. Furthermore, we also identifies several narrow fully heavy tetraquark states that may be observable in future experiments.
@@ I Introduction
Quantum Chromodynamics (QCD) allows for a variety of exotic hadronic structures beyond conventional mesons and baryons, such as multiquark states, glueballs, and hybrid states [Godfrey:1985xj, Klempt:2007cp, Lebed:2016hpi]. Among these, tetraquark states have become a major focus in hadron physics. Since the Belle Collaboration’s discovery of the $X(3872)$ resonance in 2003 [Belle:2003nnu], a growing number of exotic hadrons have been observed experimentally [ParticleDataGroup:2024cfk]. These discoveries not only confirm the existence of exotic hadronic states but also inspire further theoretical and experimental investigations into their properties. Among the tetraquark states, fully heavy tetraquarks represent a particularly intriguing class. In such systems, light-meson exchange is forbidden, and the interaction is dominated by short-range color forces mediated by gluon exchange, suggesting a compact multiquark configuration.
@@
In recent years, several candidates for fully charmed tetraquark states have been observed experimentally. In 2020, the LHCb Collaboration observed a broad structure above twice the $J/\psi$ mass threshold ranging from 6.2 to 6.8 GeV and a narrower structure at about 6.9 GeV, called $X(6900)$ [LHCb:2020bwg]. Subsequently, the CMS and ATLAS Collaborations also researched the di-$J/\psi$ channel and both of them confirmed the existence of the $X(6900)$ [CMS:2023owd, ATLAS:2023bft]. Notably, the CMS Collaboration also reported two new structures, $X(6600)$ and $X(7200)$, in the di-$J/\psi$ invariant mass spectrum [CMS:2023owd]. The ATLAS Collaboration also observed the new structures, $X(6400)$ and $X(6600)$, in the di-$J/\psi$ channel and $X(7200)$ in the $J/\psi$-$\psi(2S)$ invariant mass spectrum [ATLAS:2023bft]. Furthermore, analysis of LHCb data generating di-$J/\Psi$ in proton-proton collisions within a coupled-channel framework based on double-vector-charmonium channels indicates the existence of $X(6200)$ [Song:2024ykq]. Very recently, the CMS Collaboration determined the quantum numbers of these three fully charmed tetraquark states $X(6600)$, $X(6900)$, and $X(7100)$ to be $J^{PC}=2^{++}$ [CMS:2025fpt]. These experiments have motivated a large number of theoretical studies on fully charmed tetraquark states [Bedolla:2019zwg, Zhang:2020xtb, Wang:2022xja, Dong:2022sef, Faustov:2022mvs, Dong:2020nwy, Liang:2021fzr, Yu:2022lak, Kuang:2023vac, Liu:2021rtn, Debastiani:2017msn, Jin:2020jfc, Zhang:2020hoh, Wang:2019rdo, liu:2020eha, Lu:2020cns, Giron:2020wpx, Karliner:2020dta, universe7040094, Dong:2025coi, Dosch:2020hqm, Chen:2024orv, NgaOngodo:2025mkt]. Many theoretical studies suggest that $X(6900)$ is a $2S$-wave fully charmed tetraquark state [liu:2020eha, Lu:2020cns, Giron:2020wpx, Karliner:2020dta, universe7040094, Dong:2025coi], while others propose it as an $1S$-wave state with total spin quantum number $S=2$ [Dosch:2020hqm]. For $X(6600)$, Some works interpret it as a $2S$-wave state [Dong:2025coi, NgaOngodo:2025mkt], whereas others regard it as a $1S$-wave state [liu:2020eha]. There are also studies that suggest $X(7200)$ could be a $3S$-wave state [Zhu:2020xni, universe7040094, Dong:2025coi] or a $2P$-wave state [Yu:2022lak].
@@
In addition, some new experimental evidence for fully bottom tetraquark states has been reported. In 2016, the CMS Collaboration announced the first observation of $\Upsilon(1S)$ pair production in $pp$ collisions at $\sqrt{s}=8$ TeV [CMS:2016liw]. They found an excess with a global significance of 3.6 $\sigma$ in the $4\mu$ channel. The invariant mass of this excess is $18.4\pm 0.1\pm 0.2$ GeV [durgut2018aps]. Two years later, the LHCb Collaboration performed a research for the decay of the beautiful tetraquark, $X_{b\bar{b}b\bar{b}}$, to the $\Upsilon(1S)\mu^{+}\mu^{-}$ final state. However, no significant excess is seen for any mass hypothesis in the range [17.5, 20.0][17.5,\ 20.0] GeV [LHCb:2018uwm]. In 2019, other production of double $\Upsilon(1S)$ is observed in Cu + Au collisions at the Relativistic Heavy Ion Collider by ANDY Collaboration. They shows evidence of a significant peak at $M=18.12\pm 0.15\pm 0.6$ GeV [ANDY:2019bfn], which is compatible with the first observation of an all-$b$ tetraquark. The fully bottom tetraquark states have also been the subject of extensive theoretical investigations [Anwar:2017toa, Bai:2016int, Chen:2019dvd]. Given that only a few experiments have observed these possible fully bottom tetraquark states, further experimental investigations are required to confirm their existence.
@@
Fully heavy tetraquark states, especially the fully charmed and fully bottom tetraquark states, have attracted extensive attention from theorists. A wide range of theoretical approaches have been employed to study them, including the QCD sum rules [Chen:2016jxd, Chen:2022sbf, Wang:2022xja], a variety of quark models [Faustov:2022mvs, Liu:2019zuc, Liu:2021rtn, Jin:2020jfc, Zhang:2022qtp, Wang:2021kfv, Meng:2024yhu], the diquark models [Dong:2022sef, Debastiani:2017msn, Giron:2020wpx, Mutuk:2021hmi], the chromomagnetic model [Weng:2020jao, Deng:2020iqw, Wu:2016vtq], the diffusion Monte Carlo approach [Bai:2016int, Gordillo:2020sgc, Ma:2022vqf], and the other phenomenological models [Karliner:2020dta, universe7040094, Dong:2025coi, Dosch:2020hqm, NgaOngodo:2025mkt]. A few studies suggest that the fully heavy tetraquark states $cc\bar{c}\bar{c}$ and $bb\bar{b}\bar{b}$ could be bound states. In Ref. [Wu:2016vtq], the lowest mass of the $cc\bar{c}\bar{c}$ system lies below the di-$\eta_{c}$ threshold, while the lightest S-wave $bb\bar{b}\bar{b}$ state appears below the $\eta_{b}\eta_{b}$ threshold in Ref. [Bedolla:2019zwg, Agaev:2023wua]. However, the majority of investigations conclude that the $S$-wave fully heavy tetraquarks are located above the corresponding double-meson thresholds. For example, in Ref. [Hughes:2017xie], the authors used the first-principles lattice nonrelativistic QCD methodology to study the low-lying spectrum of the $bb\bar{b}\bar{b}$ tetraquark states with quantum numbers $0^{++}$, $1^{+-}$ and $2^{++}$. Unfortunately, they found no evidence of a QCD bound fully heavy tetraquark below the lowest non-interacting thresholds in the channels studied. It is evident that theorists hold diverse views regarding the new experimental observations, offering a variety of interpretations and even arriving at some contradictory conclusions. Meanwhile, some progress has been made in the study of mixed charm-bottom tetraquark states [Mutuk:2022nkw, Agaev:2024pej, Wu:2024hrv]. Similarly, many issues remain unresolved.
@@
Nowadays, a variety of new hadrons are continuously being observed in experiments, while theoretical progress remains relatively limited. Further theoretical studies are still required to deepen our understanding of the mechanisms underlying strong interactions. In this work, we adopted the diquark–antidiquark configuration to address the prohibition of light meson exchange in the fully heavy tetraquark state. We then systematically investigated the mass spectra and decay channels of the fully heavy tetraquark states.
@@
This paper is organized as follows. In Sec. II, we present the effective Hamiltonian employed in this study and provide the flavor-color-spin basis wave functions corresponding to different quantum numbers. Subsequently, we introduce the formulae used for calculating the strong decay widths. In Sec. III, we present the model parameters required for the calculations and describe the methods for determining these parameters. We then show the calculation results for the possible fully heavy tetraquark states and provide an analysis of their decay channels. We summarize in Sec. IV. The nature units $c=\hbar=1$ are used throughout.
@@ II.1 Effective Hamiltonian
The bound state system formed by four fully heavy quarks can be seen as a unique example of direct interaction between quarks and gluons. Consequently, the interaction in this system is mainly manifested as the one gluon exchange process [DeRujula:1975qlm, Guo:2021yws]. For the fully heavy tetraquark systems, the improved nonrelativistic effective Hamiltonian reads [Maiani:2004vq, Drenska:2008gr]
@@
$H_{SS}={}$ $\sum_{i<j}2\kappa_{ij}\left(\bm{S}_{i}\cdot\bm{S}_{j}\right),$ (2)
@@
$H_{L}={}$ $B\frac{\bm{L}\left(\bm{L}+1\right)}{2}.$
Here, $m_{i}$ and $\bm{S}_{i}$ denote the mass and spin of the $i$-th constituent quark, respectively. The parameter $\kappa_{ij}$ is related to the strong coupling constant, the effective mass of the constituent quarks, the color coupling strength, and the spatial feature of the hadron. $H_{SS}$ represents the spin-spin interaction of each quark pair in the tetraquark. $H_{SL}$ and $H_{L}$ describe spin-orbit coupling and pure orbital contribution, respectively. $\bm{S}_{\mathcal{Q}}$ in $H_{SL}$ represents the spin operator of the diquark configuration $\mathcal{Q}=[Q_{1}Q_{2}]$. The orbital angular momentum $\bm{L}$ describes the relative motion between the diquark $\mathcal{Q}$ and the anti-diquark $\bar{\mathcal{Q}}^{\prime}$. The coefficients $A$ and $B$ are phenomenological parameters to be determined from experimental data.
@@ II.2 Basis wave functions
To obtain the mass spectra and decays of the fully heavy tetraquark states, we need to determine their wave functions. Here we only show the basis wave functions of ground states and we represent them as the symbol $|{(Q_{1}Q_{2})_{s}^{c}(\bar{Q}_{3}\bar{Q}_{4})_{s}^{c}}\rangle$, where the superscript denotes the color and the subscript denotes the spin.
@@
Now, the basis wave functions of the fully heavy tetraquark system based on flavor, spin and color can be constructed as follows.
(1) For the quantum numbers $J^{P}=0^{+}$, the basis vectors $\alpha_{i}\,(i=1,2,3,4)$ can be labeled as,
@@
(2) For the axial vector tetraquark states with quantum numbers $J^{P}=1^{+}$, the basis vectors $\beta_{i}\,(i=1,2,\cdots,6)$ can be expressed as,
@@
(3) For $J^{P}=2^{+}$ states, the basis vectors $\gamma_{i}\,(i=1,2)$ of the fully heavy tetraquark states read
@@
The $\bar{\delta}_{12}$ (or $\bar{\delta}_{34}$) means that if $Q_{1}Q_{2}$ (or $\bar{Q}_{3}\bar{Q}_{4}$) have the same flavor then this state can not exist. We also consider the charge conjugation and excited states with $L=1$. The complete wave functions are listed in Appendix A.
@@ II.3 Decay width
By inserting the basis vectors into the Hamiltonian (1), we can obtain the mass spectra and corresponding wave functions of the fully heavy tetraquark states. We can then calculate the width of each tetraquark state decaying into the two meson final states.
For a particle of mass $M$ and four-momentum $P$ decaying into two particles with momenta $p_{1}$ and $p_{2}$, the partial decay rate can be written as
$d\Gamma=\frac{\left(2\pi\right)^{4}}{2M}|\mathcal{A}|^{2}d\Phi\left(P;p_{1},p_{2}\right),$ (7)
where $\mathcal{A}$ is the amplitude connecting the initial state to the final state. And the symbol $d\Phi$ is an element of 2-body phase space given by
$d\Phi\left(P;p_{1},p_{2}\right)$ $=\delta^{4}\left(P-p_{1}-p_{2}\right)$ (8)
$\times\frac{d^{3}p_{1}}{\left(2\pi\right)^{3}2E_{1}}\frac{d^{3}p_{2}}{\left(2\pi\right)^{3}2E_{2}}.$
@@
$|\bm{p}_{1}|=|\bm{p}_{2}|=\frac{1}{2M}\sqrt{\lambda\left(M^{2},m_{1}^{2},m_{2}^{2}\right)}$ (9)
where $\lambda\left(\alpha,\beta,\gamma\right)=\alpha^{2}+\beta^{2}+\gamma^{2}-2\alpha\beta-2\alpha\gamma-2\beta\gamma$ is the Källén function. The notations $m_{1}$ and $m_{2}$ represent the masses of the subparticles in the final state of the two-body decay, respectively. Thus, the partial decay rate can be written as
@@ III.1 Model parameters
To calculate the ground-state masses and corresponding wave functions of fully heavy tetraquark states, we need to determine the constituent masses of the charm and bottom quarks, as well as the parameters $\kappa_{ij}$ associated with quark pairs. Similar to the method in Ref.  [Kang:2025xqm, Liu:2025jyn], we can obtain these model parameters by fitting experimental values of conventional hadrons. For conventional hadrons that have not yet been observed experimentally, we used calculations based on lattice QCD [Alexandrou:2017xwd, Brown:2014ena]. The effective masses of the charm quark and the bottom quark can be obtained as $m_{c}=1556$ MeV and $m_{b}=4755$ MeV, respectively.
Without loss of generality, in the conventional hadron system, the color coupling between any two quarks can only be $1$ or $\bar{3}$, where $1$ corresponds to a pair of quark-antiquark forming a meson, and $\bar{3}$ corresponds to the coupling of any two quarks in a baryon. Thus we first determine $(\kappa_{Q\bar{Q}})^{1}$ with mesons and $(\kappa_{QQ})^{\bar{3}}$ with baryons.
@@
$M=m_{Q_{1}}+m_{Q_{2}}+(\kappa_{Q_{1}\bar{Q}_{2}})^{1}\left[J\left(J+1\right)-\frac{3}{2}\right].$ (12)
With these equation we can calculate the value of $(\kappa_{Q_{1}\bar{Q}_{2}})^{1}$. Their numerical results are tabulated in Table 1.
Table 1: Parameter $\kappa_{ij}$ (in MeV) is used for $Q\bar{Q}$ with a singlet color configuration (11) and $QQ$ with an anti-triplet color configuration ($\bar{3}$). Due to symmetry, these parameters satisfy $\kappa_{ij}=\kappa_{ji}$. And these parameters are derived from fitting the $1S$-wave mesons and baryons.
$(\kappa_{c\bar{c}})^{1}$ $(\kappa_{b\bar{b}})^{1}$ $(\kappa_{b\bar{c}})^{1}$ $(\kappa_{cc})^{\bar{3}}$ $(\kappa_{bb})^{\bar{3}}$ $(\kappa_{bc})^{\bar{3}}$
@@
To determine $(\kappa_{QQ})^{\bar{3}}$ with baryons, Eq. (1) can be reformulated as
$M=$ $\sum_{i=1}^{3}m_{i}+2\left(\kappa_{Q_{1}Q_{2}}\right)^{\bar{3}}\left(\bm{S}_{Q_{1}}\cdot\bm{S}_{Q_{2}}\right)+\left(\kappa_{Q_{1}Q_{3}}\right)^{\bar{3}}\left(\bm{S}_{Q_{1}}\cdot\bm{S}_{Q_{3}}\right)$
$+$ $\left(\kappa_{Q_{2}Q_{3}}\right)^{\bar{3}}\left(\bm{S}_{Q_{2}}\cdot\bm{S}_{Q_{3}}\right).$ (13)